OpenStudy (anonymous):

solve x-1=sqrt (x^4-2x^3)/(x^2-4)

OpenStudy (anonymous):

Can you take a picture of it please?

OpenStudy (anonymous):

The question

OpenStudy (anonymous):

Few moments I'll try

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

X=2/3

OpenStudy (anonymous):

I have to go to work, keep replying please :) I will be back on later tonight

OpenStudy (campbell_st):

start by factoring the part inside the square root you get x^3(x -2)/(x-2)(x+2) remove the common factor leaves x^3/(x + 2) so square both sides of the equation (x -1)^2 = x^3/(x +2) multiply both sides by (x + 2) and you have (x+2)(x -1)^2 = x^3 distribute the left had side, collect like terms and you should have a quadratic to solve hope it helps

OpenStudy (campbell_st):

after distributing you get a linear equation... so its easy to solve

OpenStudy (anonymous):

Talking about the thing under the square root, The numerator,x^4-2x^3 Take x^3 common..x^3 (x-2) And in denominator..use (a^2-b^2)=(a+b)(a-b) So denominator now is,(x+2)(x-2) now (x-2) was in numerator too,so both of them cancel out.. we are left with, x-1=square root of (x^3/x+2) Square both the sides, (x-1)^2=(x^3/x+2) (x+2)(x^2+1-2x)=x^3 x^3+x-2x^2+2x^2+2-4x=x^3 Everything cancels out infact,leaving us with, 3x=2 Therefore,x=2/3

OpenStudy (anonymous):

x-1= sqrt (x^4-2x^3)/(x^2-4) (x-1)^2 = [x^3(x-2)]/[(x-2)(x+2)] (x^2-2x+1)=x^3/(x+2) ...... * (x+2) = x^3 -3x+2=0 x=2/3

OpenStudy (anonymous):

omg, thank you guys SO much :D